# Concurrent Composition for Differentially Private Continual Mechanisms

### v2

@@ 4.3. Concurrent Composition of CMs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$
For $i\in[k]$, let $\alpha_{i}=(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})$ denote a creation query. For any input sequence of messages $m_{1},\dots,m_{t}$, the outcome of $f^{\alpha_{1},\dots,\alpha_{k}}$ returns $\top$ if and only if all of the following conditions hold:
@@
  3. (iii)
    For $i\in[k]$, the sequence of $(q^{0},q^{1})$s in the messages of the form $\left((q^{0},i),(q^{1},i)\right)$ is $f_{i}$-valid.
@@
For $1\leq i\leq k$, let $\mathcal{M}_{i}$ be a continual mechanism with initial state space $S_{\mathcal{M}_{i}}^{\text{init}}=\{s_{i}\}$ such that $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP w.r.t. a verification function $f_{i}$ against adaptive adversaries, and $\mathcal{V}[f_{i}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}_{i}$ has discrete answer distributions. For $\epsilon>0$ and $0\leq\delta\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent composition of $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is also $(\epsilon,\delta)$-DP against adaptive adversaries.
@@
Since $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}$ is an IPM with a single initial state, Lemma 2.15 implies that if ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$ is $(\epsilon,\delta)$-DP, then $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$ is also $(\epsilon,\delta)$-DP. Hence, for any adversary $\mathcal{A}$, the views of $\mathcal{A}$ when interacting with the mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((0)_{i=1}^{k}\right)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((1)_{i=1}^{k}\right)$ are $(\epsilon,\delta)$-indistinguishable. As we will show below, by construction, these views are identical to those obtained by interacting with the original mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, implying that the concurrent composition of $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is $(\epsilon,\delta)$-DP.
@@ 4.4. Concurrent Composition of CMs with Fixed Parameters
For any input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$ returns $\top$ if and only if all of the following conditions hold:
@@
  3. (iii)
    For each $i\in[k]$, the sequence of $(q^{0},q^{1})$s in the messages of the form $\left((q^{0},i),(q^{1},i)\right)$ is $f_{i}^{\prime}$-valid.
We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$-concurrent composition of CMs as the concurrent composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$. A slight adaptation of the proof of Theorem 4.10 implies the following theorem.
@@
For $\epsilon,\epsilon_{1},\dots,\epsilon_{k}>0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent composition of $k$ CMs with the privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ is also $(\epsilon,\delta)$-DP against adaptive adversaries.
@@
The proof is identical to the proof of Theorem 4.10 with the following modification in the construction of $\mathcal{P}$: Recall that $(r_{1},\dots,r_{k})$ denotes the output of ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$. In the proof of Theorem 4.10, the $i$-th creation query was known to be $(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})$. Consequently, $\mathcal{T}_{i}$ expected $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ as its right mechanism, and when $\mathcal{T}_{i}$ needed a response from its right mechanism, $\mathcal{P}$ sent $r_{i}$ to it. However, in the concurrent composition with fixed parameters, the IPM requesting the sample from $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ need not necessarily be $\mathcal{T}_{i}$. To handle this, we modify $\mathcal{P}$ so that it maps each output sample $r_{i}$ to the appropriate IPM $\mathcal{T}_{j}$ needing a sample from $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$.
Let $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ denote the $j$-th pair of identical creation queries received by $\mathcal{P}$ from the verifier, and let $\mathcal{T}_{j}$ be the corresponding IPM given by Lemma 4.6. When $\mathcal{T}_{j}$ requests a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, $\mathcal{P}$ iterates over $\sigma_{1}$ to find the smallest $i^{*}\in[k]$ such that the $i^{*}$-th element of $\sigma_{1}$ is not null and $\epsilon_{j}^{\prime}=\epsilon_{i^{*}}$, $\delta_{j}^{\prime}=\delta_{i^{*}}$. After sending $r_{i^{*}}$ to $\mathcal{T}_{j}$, $\mathcal{P}$ updates $\sigma_{1}$ and sets $r_{i}$ to null. This adjustment ensures that even when there are duplicate privacy parameters among $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$, each sample $r_{i}$ is sent to at most one $\mathcal{T}_{j}$. Note that by the definition of $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$, $i^{*}$ always exists. ∎
@@ 5. Concurrent Parallel Composition of CMs
In Section 5.1, we formally define this composition and extend the result of Theorem 4.12 to the concurrent $k$-sparse parallel composition of CMs when $\delta_{1}=\dots=\delta_{k}=0$. In Section 5.2, we provide a counter example that demonstrates why such an extension is impossible for $\delta_{j}>0$.
In Section 5.3, we restrict the verification functions $f_{i}$ of the $k$ mechanisms receiving neighboring input and we show that Theorem 4.12 holds for the concurrent $k$-sparse parallel composition of CM in this scenario without any constraints on the values of $\delta_{j}$.
5.1. Definition and Composition Theorem for Pure DP
@@
For any input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$ returns $\top$ if and only if all of the following conditions hold:
@@
  3. (iii)
    For every $j\in\mathcal{J}$, the sequence of $(q_{0},q_{1})$s in the messages of the form $\left((q_{0},j),(q_{1},j)\right)$ must be $f_{i}$-valid. Note that by definition, for every index $j\notin\mathcal{J}$ and every message of the form $\left((q_{0},j),(q_{1},j)\right)$, it holds that $q_{0}=q_{1}$.
@@
For $\epsilon,\epsilon_{1},\dots,\epsilon_{k}\geq 0$ and $0\leq\delta\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $(\epsilon,\delta)$-DP, then the concurrent $k$-sparse parallel composition of CMs with the privacy parameters $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$ is also $(\epsilon,\delta)$-DP against an adaptive adversary.
@@
The proof is identical to the proof of Theorem 4.12, except for the definition of IPMs $\mathcal{T}_{j}$ that are executed by the IPM $\mathcal{P}$. Let $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ denote the $j$-th pair of identical creation queries received by $\mathcal{P}$ from $\mathcal{V}[f]$. To ensure that the proof of Theorem 4.12 remains valid, we must guarantee that the newly defined IPMs $\mathcal{T}_{j}$ satisfy the following properties needed by the proof of Theorem 4.12. These are the only properties that this proof requires for the $\mathcal{T}_{j}$s:
  • The mechanisms $\mathcal{T}_{j}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ are equivalent, and
  • the multiset of privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ corresponding to IMPs $\mathcal{T}_{j}$ that are interacting with their right mechanism $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ is a sub-multiset of $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$, implying that the remaining $\mathcal{T}_{j}$ never interact with a $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$.
The new definition of $\mathcal{T}_{j}$ is as follows:
  • If $\delta_{j}^{\prime}>0$, then $\mathcal{T}_{j}$ acts identical to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$: $\mathcal{T}_{j}$ has a single initial state $\left((),0,s^{\text{init}}\right)$, where $s^{\text{init}}$ is the single initial state of $\mathcal{M}_{j}^{\prime}$. Upon receiving a right message, $\mathcal{T}_{j}$ halts the communication, and upon receiving a left message, it updates its state and responds identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$.
  • If $\delta_{j}^{\prime}=0$, then $\mathcal{T}_{j}$ is the IPM corresponding to $\mathcal{M}_{j}^{\prime}$ given by Lemma 4.7.
By Definition 5.1 (ii), if $\delta_{j}^{\prime}>0$, then all future messages $\left((q_{0},j),(q_{1},j)\right)$ satisfy $q_{0}=q_{1}$. Therefore, by the above definition, for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}_{j}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ are identical, and $\mathcal{T}_{j}$ never interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$.
Moreover, when $\delta_{j}^{\prime}=0$, by Definition 5.1 (iii), Lemma 4.7, and Remark 4.8, for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}_{j}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ are identical, and $\mathcal{T}_{j}$ interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ at most once after receiving a pair of non-identical queries.
In the concurrent $k$-sparse parallel composition, an unlimited number of CMs can be created, but only at most $k$ of them (here, with $\delta_{j}^{\prime}=0$) can receive non-identical query pairs. By Definition 5.1 (ii), the privacy parameters of these mechanisms is a sub-multiset of $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$. By design, the IPMs $\mathcal{T}_{j}$ corresponding to these mechanisms may require a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},0}$. ∎
5.2. Counter Example.
In this section, we illustrate that extending the result of Theorem 5.2 to general $\delta_{j}>0$ is impossible. Specifically, let $k=1$, $\epsilon_{1}=0$, and $\delta_{1}=\delta\in(0,1]$. We construct, an adversary $\mathcal{A}_{\delta}$ that interacts with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ and successfully determines the secret bit $b\in\{0,1\}$. Our construction proceeds in three steps: (1) introducing a verification function $g$, (2) constructing a CM $\mathcal{M}_{\delta}$ that is $(0,\delta)$-DP w.r.t. $g$, and (3) designing $\mathcal{A}_{\delta}$.
@@
[…]
@@
Once $\mathcal{A}_{\delta}$ receives $\bot$, without increasing $s$, it sends $\left((0,s/2),(1,s/2)\right)$. That is, $\mathcal{A}_{\delta}$ asks the pair $(0,1)$ from the last created instance of $\mathcal{M}_{\delta}$, which has returned $\bot$. If the initial state of the identifier $\mathcal{I}$ in $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $b$, then $\mathcal{M}_{\delta}$ receives the query $b$, and by definition it will return $b$ as output, revealing the secret bit $b$.

### v3

@@ 4.3. Concurrent Composition of CMs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$
For $i\in[k]$, let $\alpha_{i}=(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})$ denote a creation query. For every $t\in\mathbb{N}$ and every input sequence of messages $m_{1},\dots,m_{t}$, the verification function $f^{\alpha_{1},\dots,\alpha_{k}}$ returns $\top$ if and only if all of the following conditions hold:
@@
For $1\leq i\leq k$, $\epsilon_{1},\dots,\epsilon_{k}>0$, and $0\leq\delta_{1},\dots,\delta_{k}\leq 1$, let $\mathcal{M}_{i}$ be a continual mechanism with initial state space $S_{\mathcal{M}_{i}}^{\text{init}}=\{s_{i}\}$ such that $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP w.r.t. a verification function $f_{i}$ against adaptive adversaries, and $\mathcal{V}[f_{i}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}_{i}$ has discrete answer distributions. For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent composition of $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is also $(\epsilon,\delta)$-DP against adaptive adversaries.
@@
Since $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}$ is an IPM with a single initial state, Lemma 2.16 implies that if ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$ is $(\epsilon,\delta)$-DP, then $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$ is also $(\epsilon,\delta)$-DP. Hence, for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ when interacting with the mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((0)_{i=1}^{k}\right)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((1)_{i=1}^{k}\right)$ are $(\epsilon,\delta)$-indistinguishable. As we will show below, by construction, these views are identical to those obtained by interacting with the original mechanisms $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, implying that the concurrent composition of $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is $(\epsilon,\delta)$-DP.
@@ 4.4. Concurrent Composition of CMs with Fixed Parameters
For every $t\in\mathbb{N}$ and every input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$ returns $\top$ if and only if all of the following conditions hold:
@@
We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$-concurrent composition of CMs as the concurrent composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$. A slight adaptation of the proof of Theorem 4.9 implies the following theorem. For convenience, in the rest of the paper, we represent a pair of identical creation queries $(\mathcal{M},s,\epsilon,\delta,f)$ by $(\mathcal{M},s,\epsilon,\delta,f)^{2}$.
@@
For $\epsilon_{1},\dots,\epsilon_{k}>0$, $\epsilon\geq 0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent composition of $k$ CMs with the privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ is also $(\epsilon,\delta)$-DP against adaptive adversaries.
@@
The proof is identical to the proof of Theorem 4.9 with a modification in the construction of $\mathcal{P}$: Recall that $(r_{1},\dots,r_{k})$ denotes the output of ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$. In the proof of Theorem 4.9, the $i$-th creation query was known in advance to be $(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})$. Consequently, $\mathcal{T}_{i}$ expected $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ as its right mechanism, and when $\mathcal{T}_{i}$ needed a response from its right mechanism, $\mathcal{P}$ sent $r_{i}$ to it. However, in the concurrent composition with fixed parameters, the IPM requesting the sample from $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ need not necessarily be the $i$-th IPM, $\mathcal{T}_{i}$. To handle this, we modify $\mathcal{P}$ so that it maps each output sample $r_{i}$ to the appropriate IPM $\mathcal{T}_{j}$ needing a sample from $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$. The implementation details are as follows.
Let $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ denote the $j$-th pair of identical creation queries received by $\mathcal{P}$ from the verifier, and let $\mathcal{T}_{j}$ be the corresponding IPM given by Lemma 4.6. In the proof of Theorem 4.9, the state of $\mathcal{P}$ consists of a pair of two sequences $\sigma_{1}=(r_{1},\dots,r_{k})$ and $\sigma_{2}=\left((\mathcal{T}_{j},t_{j})\right)_{j=1}^{k}$, where $r_{j}$ is the randomized response defined above and $t_{j}$ denotes the current state of $\mathcal{T}_{j}$. We extend the state of $\mathcal{P}$ to additionally store an index $y_{i}$ with each $r_{i}$ in $\sigma_{1}$ and the privacy parameters $\epsilon_{j}^{\prime},\delta_{j}^{\prime}$ with each $\mathcal{T}_{j}$ in $\sigma_{2}$. Variable $y_{i}$ is initialized to 0 and used to record the index $j$ of the mechanism $\mathcal{T}_{j}$ associated with $r_{i}$
Upon receiving $(r_{1},\dots,r_{k})$ from ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]$, $\mathcal{P}$ sets $\sigma_{1}=\left((r_{1},y_{1}=0),\dots,(r_{k},y_{k}=0)\right)$. Also, upon receiving a pair of creation queries $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$, $\mathcal{P}$ appends the tuple $(\mathcal{T}_{j},t_{j},\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ to $\sigma_{2}$. When $\mathcal{T}_{j}$ requests a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, $\mathcal{P}$ iterates over $\sigma_{1}$ to find the smallest $i^{*}\in[k]$ such that $y_{i^{*}}=0$, $\epsilon_{j}^{\prime}=\epsilon_{i^{*}}$, and $\delta_{j}^{\prime}=\delta_{i^{*}}$. It then assigns $y_{i^{*}}=j$ and sends $r_{i^{*}}$ to $\mathcal{T}_{j}$ as a right response. This adjustment ensures that even when there are duplicate privacy parameters among $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$, each sample $r_{i}$ is sent to at most one $\mathcal{T}_{j}$. Note that by the definition of $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$, $i^{*}$ always exists. ∎
@@ 5. Concurrent Parallel Composition of CMs
In Section 5.1, we formally define this composition and show that the result of Theorem 4.11 cannot be extended to the concurrent $k$-sparse parallel composition of CMs when any $\delta_{j}>0$. In Section 5.2, we prove a parallel composition theorem that matches the privacy guarantee of Theorem 4.11 when only purely differentially privacy CMs are created. Finally, in Section 5.3, we consider a restricted class of verification functions $f_{i}$ for the $k$ mechanisms receiving neighboring inputs. Under certain assumptions about these functions, we show that the privacy guarantee of Theorem 4.11 still holds even if the created mechanism are approximately dp.
5.1. Definition and Counterexample
@@
For every $t\in\mathbb{N}$ and every input sequence of messages $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty}$ returns $\top$ if and only if all of the following conditions hold:
@@
The following theorem shows that for $k=1$, $\epsilon_{1}=0$, and any $\delta_{1}\in(0,1]$, the $f^{0,\delta_{1}}_{\infty}$-concurrent composition of CMs is not differentially private:
@@
For every $\delta\in(0,1]$, there exists an adversary $\mathcal{A}_{\delta}$ such that the views of $\mathcal{A}_{\delta}$ interacting with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ have disjoint supports.
@@
We construct $\mathcal{A}_{\delta}$ in three steps: (1) introducing a verification function $g$, (2) constructing a CM $\mathcal{M}_{\delta}$ that is $(0,\delta)$-DP w.r.t. $g$, and (3) designing $\mathcal{A}_{\delta}$. We then prove that for every $b\in\{0,1\}$, with probability $1$, $\mathcal{A}_{\delta}$ can determine the secret initial state $b$ by interacting with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ finitely many times. Specifically, we show that with probability $1$, the view of $\mathcal{A}_{\delta}$ interacting with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ is finite and includes $b$ as the last message. This immediately implies that the views of $\mathcal{A}_{\delta}$ interacting with $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ have disjoint supports.
@@
Step 2 (Constructing $\mathcal{M}_{\delta}$): The CM $\mathcal{M}_{\delta}:\{\bot,\top\}\times\{0,1\}\to\{\bot,\top\}\times\{0,1,\top,\bot\}$ with the initial state space $\{\top\}$ is defined as follows:
@@
Once $\mathcal{A}_{\delta}$ receives $\bot$, without increasing $s$, it sends $\left((0,s/2),(1,s/2)\right)$. That is, $\mathcal{A}_{\delta}$ asks the pair $(0,1)$ from the last created instance of $\mathcal{M}_{\delta}$, which has returned $\bot$. If the initial state of the identifier $\mathcal{I}$ in $\mathcal{V}[f^{0,\delta}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $b$, then $\mathcal{M}_{\delta}$ receives the query $b$, and by definition it will return $b$ as output, revealing the secret bit $b$. Upon receiving the response $b$, $\mathcal{A}_{\delta}$ halts the communication.